OpenStudy (anonymous):

what is tan of 2pi/3 also what is the csc of pi/4 (sorry still a bit confused)

OpenStudy (psymon):

Do you have a unit circle handy on you to help ya out with this?

OpenStudy (anonymous):

yes im looking at it. but I lost the one that has all my notes :(

OpenStudy (psymon):

Ah x_x Well, let's start out with knowing that tangent is the same and sin divided by cosine. SO if we want tangent of 2pi/3, we take the sin of 2pi/3 and divide it by the cos 2pi/3. If you can look up those values then we can find out what that is : )

OpenStudy (psymon):

\[\frac{ \sin(\frac{ 2\pi }{ 3 }) }{ \cos (\frac{ 2\pi }{ 3 }) }\]

OpenStudy (anonymous):

sin of 2pi/3 is \[\sqrt{3}\div2\] and the cos is -1/2

OpenStudy (psymon):

Correct. so now divide those 2: \[\frac{ \frac{ \sqrt{3} }{ 2 } }{ \frac{ -1 }{ 2 } } \]

OpenStudy (psymon):

Or simply: \[\frac{ \sqrt{3} }{ 2 }*-2 \]

OpenStudy (anonymous):

so would it be -sqrt of 6?

OpenStudy (psymon):

Nope. Just cancel out the two on top and bottom in what I have posted above.

OpenStudy (psymon):